Statistics problem #10
IQ scores are designed to have a mean of 100 and a standard deviation of 15.
People with IQ scores of 130+ are eligible for joining MENSA.
What percentage of people are eligible for MENSA?
What percentage of people have IQ scores between 90 and 105?
An individual has an IQ of 86. What is his/her percentile?
An individual falls at the 53rd percentile. What is his/her IQ Score?

Answer by Boreal(15235)

z=(x-mean)/sd
a. (130-100)/15=2
P(z =2 or more) is 0.0228
----------------
this is (90-100)/15 or -2/3 to (105-100)/15 or +1/3
probability z is between these is 0.3781
---------------
=(86-100)/15=-14/15, which is 0.1753 or 18th percentile.
--------------
want z for probability of 0.53
that is +0.08
0.08=(x-mean)/15
1.20=x-100
x=101.2